OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (zzr0ck3r):

what is the slope?

OpenStudy (zzr0ck3r):

(10-1)/(2-(-1))

OpenStudy (zzr0ck3r):

@shutupron ?

OpenStudy (anonymous):

9 and 1

OpenStudy (anonymous):

there isnt a slope?

OpenStudy (zzr0ck3r):

slope = (10-1)/(2-(-1)) = ?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

?

OpenStudy (zzr0ck3r):

what is 10-1?

OpenStudy (anonymous):

9

OpenStudy (zzr0ck3r):

what is 2-(-1)?

OpenStudy (anonymous):

3

OpenStudy (zzr0ck3r):

what is 9/3?

OpenStudy (anonymous):

3

OpenStudy (zzr0ck3r):

ok slope = 3

OpenStudy (zzr0ck3r):

so y=mx+b = 3x+b now we need to find b use the point (1,10) x=1 y = 10 10=3(1)+b what is b?

OpenStudy (anonymous):

3

OpenStudy (zzr0ck3r):

10=3+b what is b?

OpenStudy (zzr0ck3r):

you might need to back track and learn some basics.

OpenStudy (zzr0ck3r):

10=3+b

OpenStudy (zzr0ck3r):

what is b?

OpenStudy (zzr0ck3r):

if you have 3 apples and I give you some more apples and you now have 10 apples, how many apples did I give you? @shutupron

OpenStudy (anonymous):

7

OpenStudy (zzr0ck3r):

yes, so your equation is y=3x+7

OpenStudy (anonymous):

so a?

OpenStudy (anonymous):

3x-y=5

OpenStudy (anonymous):

I said it was A because thats the closest one to your answer !

OpenStudy (anonymous):

it says -7 not 7

OpenStudy (zzr0ck3r):

sorry -7 lol yes a

OpenStudy (anonymous):

thank you hahah